# Evidence for itinerant electron-local moment interaction in Li-doped -MnTe Tingjun Zhang

### v1

@@ Abstract
The addition of charge carriers to -MnTe may, however, impact the nature of the magnetic exchange via coupling between itinerant electrons and local moments through RKKY (indirect exchange) interactions and magnon-electron scattering processes.
These may shorten the magnon lifetime and therefore explain the broadened linewidth of spin waves throughout the BZ. The absence of a pronounced -dependence in the linewidth broadening [Figs. 3(h,i), S2] further supports this interpretation in the low-carrier density limit [sapkota_signatures_2020, hu_coupling_2023].
@@
The fact that small amounts of Li doping can dramatically affect the spin gap and the spin waves bandwidth is also supportive of the important role of itinerant electrons in this system.
The sensitive dependence of the AHE on the charge gap in undoped -MnTe [liu_StraintunableAnomalousHall_2025] further underscores the notion that the physics of this system is best described as a network of local moments that interact with itinerant charge carriers.
@@
We further test this notion by calculating the local dynamic susceptibility from the total spectral weight of our INS measurements [Fig. 4(d)], which, when integrated across all energy transfers, yields a quantitative measure of the local dynamic moment [10.1063/1.4818323, RevModPhys.87.855].
This yields a dynamic moment of 1.53(1) [1.84(6) ] for
undoped -MnTe [Li-doped -MnTe] which, when combined with the corresponding static ordered moment following the total moment sum rule (Eqn. S2), indicates a total effective moment of 5.88(4) [6.1(2) ].
These values are close to the expected for () Mn2+ ions, therefore strongly suggesting a local moment picture for altermagnetism in both compounds, while the addition of itinerant electrons serves to modify interactions between the Mn2+ spins.
In comparing Li-doped and undoped -MnTe, we find well-defined spin waves throughout the BZ with broadened energy width induced by Li doping. This is reminiscent of the broadened spin waves in ferromagnetic CrGeTe3 due to dynamical spin-lattice coupling arising from the zero-temperature lattice vibration [WOS:000895384200001]. Since spin-lattice coupling should be similar for undoped and Li-doped -MnTe, the observed spin wave broadening and damping in Li-doped -MnTe must arise from Li-doping-induced itinerant electron-local moment interactions which must be considered to understand the physics of altermagnetic candidate materials.
I Acknowledgments
@@
We thank Kirill Belashchenko for helpful discussions.
@@
The single-crystal synthesis and neutron scattering work at Rice were supported by the U.S. DOE, BES under Grant Nos. DE-SC0012311 (P.D.) and DE-SC0026179 (P.D., E.M.). Part of the materials characterization efforts at Rice is supported by the Robert A. Welch Foundation Grant No. C-1839 (P.D.). X.W. acknowledges research sponsored by the Laboratory Directed Research and Development Program of ORNL, managed by UT-Battelle, LLC, for the U.S. DOE. Part of the materials characterization efforts at University of Houston is supported by the T. L. L. Temple Foundation; the John J and Rebecca Moores Endowment; the State of Texas through the Texas Center for Superconductivity at the University of Houston; and the Robert A. Welch Foundation (00730-5021-H0452-B0001-G0512489).
@@
Supplemental Information
II Experimental Details
II.1 Sample synthesis and characterization
@@
Single crystals of -MnTe and Li-doped -MnTe were synthesized using Te self-flux method following the procedures as described in our previous work [liu_StraintunableAnomalousHall_2025]. Magnetic susceptibility measurements were performed on a single-crystal Li-doped -MnTe sample with a mass of 90.4 mg, using a Quantum Design Magnetic Property Measurement System (MPMS) equipped with a superconducting quantum interference device (SQUID) magnetometer. The sample was mounted on the sample holder using GE varnish. Both field-cooled (FC) and zero-field-cooled (ZFC) measurements were conducted with the magnetic field applied along the -axis and within the -plane.
II.2 Neutron diffraction
@@
Neutron diffraction measurements were carried out using the single crystal diffractometer TOPAZ at the Spallation Neutron Source (SNS), Oak Ridge National Laboratory (ORNL) [coates2018suite]. The experiments employed a broad neutron wavelength range of 0.4–3.5 Å. Single crystals of Li-doped -MnTe ( 100 mg) and pure -MnTe ( 50 mg) were used. The samples were mounted on the ambient goniometer, and data were collected using time-of-flight (TOF) techniques, which provided wavelength-resolved Laue diffraction patterns. Diffraction datasets for both samples were obtained at 90 K and 340 K. A nitrogen cryostream was utilized for temperature control.
@@
The data collection strategy was determined using the planning tool in NeuXtalViz [NeuXtalViz]. A multiresolution machine learning method was employed for peak integration to obtain peak intensities in three-dimensional space [reshniak2024hierarchical]. Data reduction and normalization, including the Lorentz factor, neutron TOF spectrum, and detector efficiency corrections, were carried out following the procedure reported previously [schultz2014integration]. The reduced data were saved in SHELX HKLF2 format [sheldrick2015crystal] with the neutron wavelength recorded separately for each reflection. The neutron crystal structures were solved in JANA2020 [petvrivcek2023jana2020] and refined successfully to convergence.
@@
II.3 Inelastic neutron scattering
Inelastic neutron scattering (INS) measurements on undoped -MnTe and preliminary measurements on Li-doped -MnTe were performed using the fine-resolution Fermi chopper spectrometer SEQUOIA at the Spallation Neutron Source (SNS), Oak Ridge National Laboratory (ORNL) [granroth2006sequoia, granroth2010sequoia]. For the pure -MnTe sample, a single crystal with a mass of approximately 0.46 g was mounted in the [, 0, ] scattering plane. Data were collected at K using an incident energy of meV with the high-flux chopper operating at a frequency of Hz. Measurements were performed over a total rotation range of 120 degrees about the vertical axis of the sample, with a step size of 0.5 degrees. During the experiment, the SNS accelerator operated at a beam power of 1.8 MW, and data were collected for an accumulated proton charge of approximately 0.4 Coulombs per angle step, corresponding to a counting time of roughly 5 minutes. The data obtained from SEQUOIA are presented in Figs. 2(b)–(e), 3(a)(c)(f)–(i), 4, and S2.

### v2

@@ Abstract
To verify this, we performed ARPES measurements on Li-MnTe to interrogate the behavior of the electronic band splitting in the presence of Li dopants. As with MnTe, the Fermi surface at consists of six petal-like pockets around [Fig. 4(e)] [37, 51]. Along -K, where altermagnetic splitting from SOC has previously been predicted and observed in MnTe, we find splitting of approximately the same magnitude (300 meV) [37, 51, 18], verifying the robust nature of altermagnetism in 3% Li-doped MnTe [Fig. 4(f)]. Within the scope of these experiments, no obvious changes in the band structure near the Fermi level were observed when compared to MnTe. Thus, the observed magnon damping must arise from subtle changes in the electronic structure that impact the nature of the magnetic exchange and spin dynamics via coupling between itinerant electrons and local moments through RKKY (indirect exchange) interactions and magnon-electron scattering processes, respectively.
@@
The fact that small amounts of Li doping can dramatically affect the spin gap and the spin wave bandwidth is also supportive of the important role of itinerant electrons in this system.
@@
Further, the sensitive dependence of the AHE on the charge gap in undoped MnTe [44] underscores the notion that the physics of this system can largely be described as a network of local moments that interact with itinerant charge carriers. We further test this notion by calculating the local dynamic susceptibility from the total spectral weight of our INS measurements [Fig. 4(d)], which, when integrated across all energy transfers, yields a quantitative measure of the local dynamic moment [72, 16]. Our analysis yields 1.53(1) [1.84(6) ] for pristine MnTe [Li-MnTe]. When combined with the corresponding static ordered moment (Eqn. S2), this indicates a total effective moment of 5.88(4) [6.1(2) ].
The small but finite magnon broadening is not expected to remove spectral weight, and none of the spectral weight is pushed to energies outside the measurement range. Thus, these values are close to the expected for () Mn2+ ions, therefore strongly suggesting a local moment picture for altermagnetism in both compounds despite partially itinerant character at high energies upon 3% Li doping.
In summary, magnon damping is observed in Li-MnTe along multiple high-symmetry paths near the BZ boundary at high energy, suggesting a mechanism based in subtle electronic structure changes with a characteristic energy distinct from the strongly momentum-dependent altermagnetic splitting.
Additional chirality-dependent magnon decay processes have been proposed in altermagnets [15, 13, 21], although whether these mechanisms can be directly observed by INS remains an open question.
This finding underscores the important role of itinerant electron-local moment interactions, which must be considered to understand the physics of altermagnetic candidate materials and their potential for applications in next-generation spintronic devices.
@@
We thank Kirill Belashchenko, Erick Lawrence, and Yichen Zhang for helpful discussions.
@@
The single-crystal synthesis and neutron scattering work at Rice were supported by the U.S. DOE, BES under Grant Nos. DE-SC0026179 (P.D., M.Y., and E.M.) and DE-SC0012311 (P.D.). Part of the materials characterization efforts at Rice is supported by the Robert A. Welch Foundation Grant No. C-1839 (P.D.). X.W. acknowledges research sponsored by the Laboratory Directed Research and Development Program of ORNL, managed by UT-Battelle, LLC, for the U.S. DOE. Part of the materials characterization efforts at University of Houston is supported by the T. L. L. Temple Foundation; the John J and Rebecca Moores Endowment; the State of Texas through the Texas Center for Superconductivity at the University of Houston; and the Robert A. Welch Foundation (00730-5021-H0452-B0001-G0512489). This research was supported in part by an appointment to the ORNL GRO Program, sponsored by the U.S. DOE and administered by the Oak Ridge Institute for Science and Education.
@@
[2]O. J. Amin, A. Dal Din, E. Golias, Y. Niu, A. Zakharov, S. C. Fromage, C. J. B. Fields, S. L. Heywood, R. B. Cousins, F. Maccherozzi, J. Krempaský, J. H. Dil, D. Kriegner, B. Kiraly, R. P. Campion, A. W. Rushforth, K. W. Edmonds, S. S. Dhesi, L. Šmejkal, T. Jungwirth, and P. Wadley (2024-12)Nanoscale imaging and control of altermagnetism in MnTe.
Nature636 (8042), pp. 348–353.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[3]O. Arnold, J. Bilheux, J. Borreguero, A. Buts, S. I. Campbell, L. Chapon, M. Doucet, N. Draper, R. F. Leal, M. Gigg, et al. (2014)Mantid—Data analysis and visualization package for neutron scattering and SR experiments.
Nucl. Instrum. Methods Phys. Res. A764, pp. 156–166.
Cited by: §I.3.
[4]A. Barman, G. Gubbiotti, S. Ladak, A. O. Adeyeye, M. Krawczyk, J. Gräfe, C. Adelmann, S. Cotofana, A. Naeemi, V. I. Vasyuchka, B. Hillebrands, S. A. Nikitov, H. Yu, D. Grundler, A. V. Sadovnikov, A. A. Grachev, S. E. Sheshukova, J. Duquesne, M. Marangolo, G. Csaba, W. Porod, V. E. Demidov, S. Urazhdin, S. O. Demokritov, E. Albisetti, D. Petti, R. Bertacco, H. Schultheiss, V. V. Kruglyak, V. D. Poimanov, S. Sahoo, J. Sinha, H. Yang, M. Münzenberg, T. Moriyama, S. Mizukami, P. Landeros, R. A. Gallardo, G. Carlotti, J. Kim, R. L. Stamps, R. E. Camley, B. Rana, Y. Otani, W. Yu, T. Yu, G. E. W. Bauer, C. Back, G. S. Uhrig, O. V. Dobrovolskiy, B. Budinska, H. Qin, S. van Dijken, A. V. Chumak, A. Khitun, D. E. Nikonov, I. A. Young, B. W. Zingsem, and M. Winklhofer (2021-08)The 2021 Magnonics Roadmap.
J. Phys. Condens. Matter33 (41), pp. 413001.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[5]S. P. Bayrakci, D. A. Tennant, Ph. Leininger, T. Keller, M. C. R. Gibson, S. D. Wilson, R. J. Birgeneau, and B. Keimer (2013-07)Lifetimes of Antiferromagnetic Magnons in Two and Three Dimensions: Experiment, Theory, and Numerics.
Phys. Rev. Lett.111, pp. 017204.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[11]L. Chen, C. Mao, J. Chung, M. B. Stone, A. I. Kolesnikov, X. Wang, N. Murai, B. Gao, O. Delaire, and P. Dai (2022-JUL 12)Anisotropic magnon damping by zero-temperature quantum fluctuations in ferromagnetic CrGeTe3.
Nat. Commun.13 (1), pp. 4037.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[14]L. Coates, H. Cao, B. C. Chakoumakos, M. D. Frontzek, C. Hoffmann, A. Y. Kovalevsky, Y. Liu, F. Meilleur, A. M. dos Santos, D. A. Myles, et al. (2018)A suite-level review of the neutron single-crystal diffraction instruments at oak ridge national laboratory.
Rev. Sci. Instrum.89 (9).
External Links: LinkCited by: §I.2.
[19]J. Dzian, P. Kubaščík, S. Tázlarů, M. Białek, M. Šindler, F. Le Mardelé, C. Kadlec, F. Kadlec, M. Gryglas-Borysiewicz, K. P. Kluczyk, A. Mycielski, P. Skupiński, J. Hejtmánek, R. Tesař, J. Železný, A.-L. Barra, C. Faugeras, J. Volný, K. Uhlířová, L. Nádvorník, M. Veis, K. Výborný, and M. Orlita (2025-07)Antiferromagnetic resonance in -MnTe.
Phys. Rev. B112 (2), pp. 024433.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[21]R. Eto, M. Gohlke, J. Sinova, M. Mochizuki, A. L. Chernyshev, and A. Mook (2025)Spontaneous magnon decays from nonrelativistic time-reversal symmetry breaking in altermagnets.
Phys. Rev. B112 (9), pp. 094442.
Cited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[22]R. Ewings, A. Buts, M. Le, J. van Duijn, I. Bustinduy, and T. Perring (2016)Horace: software for the analysis of data from single crystal spectroscopy experiments at time-of-flight neutron instruments.
Nucl. Instrum. Methods Phys. Res. A834, pp. 132–142.
Cited by: §I.3.
[23]R. Ewings, J. Stewart, T. Perring, R. Bewley, M. Le, D. Raspino, D. Pooley, G. Škoro, S. Waller, D. Zacek, et al. (2019)Upgrade to the MAPS neutron time-of-flight chopper spectrometer.
Rev. Sci. Instrum.90 (3).
Cited by: §I.3.
[25]B. Flebus, D. Grundler, B. Rana, Y. Otani, I. Barsukov, A. Barman, G. Gubbiotti, P. Landeros, J. Akerman, U. Ebels, P. Pirro, V. E. Demidov, K. Schultheiss, G. Csaba, Q. Wang, F. Ciubotaru, D. E. Nikonov, P. Che, R. Hertel, T. Ono, D. Afanasiev, J. Mentink, T. Rasing, B. Hillebrands, S. V. Kusminskiy, W. Zhang, C. R. Du, A. Finco, T. van der Sar, Y. K. Luo, Y. Shiota, J. Sklenar, T. Yu, and J. Rao (2024-06)The 2024 magnonics roadmap.
J. Phys. Condens. Matter36 (36), pp. 363501.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[27]R. D. Gonzalez Betancourt, J. Zubáč, R. Gonzalez-Hernandez, K. Geishendorf, Z. Šobáň, G. Springholz, K. Olejník, L. Šmejkal, J. Sinova, T. Jungwirth, S. T. B. Goennenwein, A. Thomas, H. Reichlová, J. Železný, and D. Kriegner (2023-01)Spontaneous Anomalous Hall Effect Arising from an Unconventional Compensated Magnetic Phase in a Semiconductor.
Phys. Rev. Lett.130 (3), pp. 036702.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[28]R. D. Gonzalez Betancourt, J. Zubáč, K. Geishendorf, P. Ritzinger, B. Růžičková, T. Kotte, J. Železný, K. Olejník, G. Springholz, B. Büchner, A. Thomas, K. Výborný, T. Jungwirth, H. Reichlová, and D. Kriegner (2024-08)Anisotropic magnetoresistance in altermagnetic MnTe.
npj Spintronics2 (1), pp. 45.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[29]G. E. Granroth, D. H. Vandergriff, and S. E. Nagler (2006)SEQUOIA: A fine resolution chopper spectrometer at the SNS.
Physica B385, pp. 1104–1106.
Cited by: §I.3.
[30]G. Granroth, A. Kolesnikov, T. Sherline, J. Clancy, K. Ross, J. Ruff, B. Gaulin, and S. Nagler (2010)SEQUOIA: a newly operating chopper spectrometer at the SNS.
In J. Phys.: Conf. Ser.,
Vol. 251, pp. 012058.
Cited by: §I.3.
[31]A. Hariki, A. Dal Din, O. J. Amin, T. Yamaguchi, A. Badura, D. Kriegner, K. W. Edmonds, R. P. Campion, P. Wadley, D. Backes, L. S. I. Veiga, S. S. Dhesi, G. Springholz, L. Šmejkal, K. Výborný, T. Jungwirth, and J. Kuneš (2024-04)X-Ray Magnetic Circular Dichroism in Altermagnetic -MnTe.
Phys. Rev. Lett.132 (17), pp. 176701.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[34]X. Hu, A. Sapkota, Z. Hu, A. T. Savici, A. I. Kolesnikov, J. M. Tranquada, C. Petrovic, and I. A. Zaliznyak (2023-05)Coupling of magnetism and Dirac fermions in ${\mathrm{YbMnSb}}_{2}$.
Phys. Rev. B107 (20), pp. L201117.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[36]K. P. Kluczyk, K. Gas, M. J. Grzybowski, P. Skupiński, M. A. Borysiewicz, T. Fąs, J. Suffczyński, J. Z. Domagala, K. Grasza, A. Mycielski, M. Baj, K. H. Ahn, K. Výborný, M. Sawicki, and M. Gryglas-Borysiewicz (2024-10)Coexistence of anomalous Hall effect and weak magnetization in a nominally collinear antiferromagnet MnTe.
Phys. Rev. B110 (15), pp. 155201.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[37]J. Krempaský, L. Šmejkal, S. D’Souza, H. M, S. G, K. Uhlířová, F. Alarab, P. Constantinou, V. Strocov, D. Usanov, W. Pudelko, R. González-Hernández, A. Birk Hellenes, Z. Jansa, H. Reichlová, Z. Šobáň, R. Gonzalez Betancourt, P. Wadley, J. Sinova, D. Kriegner, J. Minár, J. Dil, and T. Jungwirth (2024-02)Altermagnetic lifting of kramers spin degeneracy.
Nature626 (7999), pp. 517–522.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[38]D. Kriegner, H. Reichlova, J. Grenzer, W. Schmidt, E. Ressouche, J. Godinho, T. Wagner, S. Y. Martin, A. B. Shick, V. V. Volobuev, G. Springholz, V. Holý, J. Wunderlich, T. Jungwirth, and K. Výborný (2017-12)Magnetic anisotropy in antiferromagnetic hexagonal MnTe.
Phys. Rev. B96 (21), pp. 214418.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[39]D. Kriegner, H. Reichlova, J. Grenzer, W. Schmidt, E. Ressouche, J. Godinho, T. Wagner, S. Y. Martin, A. B. Shick, V. V. Volobuev, G. Springholz, V. Holý, J. Wunderlich, T. Jungwirth, and K. Výborný (2017)Magnetic anisotropy in antiferromagnetic hexagonal MnTe.
Phys. Rev. B96 (21), pp. 214418.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[40]D. Kriegner, K. Výborný, K. Olejník, H. Reichlová, V. Novák, X. Marti, J. Gazquez, V. Saidl, P. Němec, V. V. Volobuev, G. Springholz, V. Holý, and T. Jungwirth (2016-06)Multiple-stable anisotropic magnetoresistance memory in antiferromagnetic MnTe.
Nat. Commun.7 (1), pp. 11623.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[42]J. Lamsal, G. S. Tucker, T. W. Heitmann, A. Kreyssig, A. Jesche, A. Pandey, W. Tian, R. J. McQueeney, D. C. Johnston, and A. I. Goldman (2013)Persistence of local-moment antiferromagnetic order in Ba${}_{1\ensuremath{-}x}$K${}_{x}$Mn${}_{2}$As${}_{2}$.
Phys. Rev. B87 (14), pp. 144418.
External Links: DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[47]K. Momma and F. Izumi (2011)VESTA 3 for three-dimensional visualization of crystal, volumetric and morphology data.
J. Appl. Crystallogr.44 (6), pp. 1272–1276.
Cited by: §I.2.
[48]D. H. Moseley, K. M. Taddei, J. Yan, M. A. McGuire, S. Calder, M. M. H. Polash, D. Vashaee, X. Zhang, H. Zhao, D. S. Parker, R. S. Fishman, and R. P. Hermann (2022-01)Giant doping response of magnetic anisotropy in MnTe.
Phys. Rev. Mater.6 (1), pp. 014404.
External Links: DocumentCited by: §II.1,
Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[53]V. Petříček, L. Palatinus, J. Plášil, and M. Dušek (2023)Jana2020–a new version of the crystallographic computing system Jana.
Z. Kristallogr. Cryst. Mater.238 (7-8), pp. 271–282.
Cited by: §I.2.
[54]V. Reshniak, X. Wang, G. Zhang, S. Liu, and J. Yin (2024)Hierarchical bayesian approach for adaptive integration of bragg peaks in time-of-flight neutron scattering data.
arXiv:2406.05133.
External Links: LinkCited by: §I.2.
[55]A. Sapkota, L. Classen, M. B. Stone, A. T. Savici, V. O. Garlea, A. Wang, J. M. Tranquada, C. Petrovic, and I. A. Zaliznyak (2020-01)Signatures of coupling between spin waves and Dirac fermions in ${\mathrm{YbMnBi}}_{2}$.
Phys. Rev. B101 (4), pp. 041111.
External Links: Link,
DocumentCited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[56]A. J. Schultz, M. R. V. Jørgensen, X. Wang, R. L. Mikkelson, D. J. Mikkelson, V. E. Lynch, P. F. Peterson, M. L. Green, and C. M. Hoffmann (2014)Integration of neutron time-of-flight single-crystal bragg peaks in reciprocal space.
J. Appl. Crystallogr.47 (3), pp. 915–921.
External Links: LinkCited by: §I.2.
[58]G. M. Sheldrick (2015)Crystal structure refinement with shelxl.
Acta Crystallogr. C71 (1), pp. 3–8.
External Links: LinkCited by: §I.2.
[59]A. K. Singh, N. Heinsdorf, A. A. Mancilla, J. Bannies, A. Maity, A. I. Kolesnikov, M. Matsuda, M. B. Stone, M. Franz, J. Gaudet, et al. (2025)Chiral spin-split magnons in the metallic altermagnet crsb.
arXiv:2511.16086.
Cited by: Evidence for itinerant electron-local moment interaction in Li-doped -MnTe.
[70]S. Toth and B. Lake (2015)Linear spin wave theory for single- incommensurate magnetic structures.
J. Phys. Condens. Matter27 (16), pp. 166002.
Cited by: §I.5.
@@
Single crystals of -MnTe and Li-doped -MnTe were synthesized using Te self-flux method following the procedures as described in our previous work [44]. Magnetic susceptibility measurements were performed on a single-crystal Li-doped -MnTe sample with a mass of 90.4 mg, using a Quantum Design Magnetic Property Measurement System (MPMS) equipped with a superconducting quantum interference device (SQUID) magnetometer. The sample was mounted on the sample holder using GE varnish. Both field-cooled (FC) and zero-field-cooled (ZFC) measurements were conducted with the magnetic field applied along the -axis and within the -plane.
I.2 Neutron diffraction
@@
Neutron diffraction measurements were carried out using the single crystal diffractometer TOPAZ at the Spallation Neutron Source (SNS), Oak Ridge National Laboratory (ORNL) [14]. The experiments employed a broad neutron wavelength range of 0.4–3.5 Å. Single crystals of Li-doped -MnTe ( 100 mg) and pure -MnTe ( 50 mg) were used. The samples were mounted on the ambient goniometer, and data were collected using time-of-flight (TOF) techniques, which provided wavelength-resolved Laue diffraction patterns. Diffraction datasets for both samples were obtained at 90 K and 340 K. A nitrogen cryostream was utilized for temperature control.
@@
The data collection strategy was determined using the planning tool in NeuXtalViz [49]. A multiresolution machine learning method was employed for peak integration to obtain peak intensities in three-dimensional space [54]. Data reduction and normalization, including the Lorentz factor, neutron TOF spectrum, and detector efficiency corrections, were carried out following the procedure reported previously [56]. The reduced data were saved in SHELX HKLF2 format [58] with the neutron wavelength recorded separately for each reflection. The neutron crystal structures were solved in JANA2020 [53] and refined successfully to convergence. Crystal structures presented in the main text were visualized using the VESTA software package [47].
@@
I.3 Inelastic neutron scattering
Inelastic neutron scattering (INS) measurements on undoped -MnTe and preliminary measurements on Li-doped -MnTe were performed using the fine-resolution Fermi chopper spectrometer SEQUOIA at the Spallation Neutron Source (SNS), Oak Ridge National Laboratory (ORNL) [29, 30]. For the pure -MnTe sample, a single crystal with a mass of approximately 0.46 g was mounted in the [, 0, ] scattering plane. Data were collected at K using an incident energy of meV with the high-flux chopper operating at a frequency of Hz. Measurements were performed over a total rotation range of 120 degrees about the vertical axis of the sample, with a step size of 0.5 degrees. During the experiment, the SNS accelerator operated at a beam power of 1.8 MW, and data were collected for an accumulated proton charge of approximately 0.4 Coulombs per angle step, corresponding to a counting time of roughly 5 minutes. The data obtained from SEQUOIA are presented in Figs. 2(b)–(e), 3(a)(c)(f)–(i), 4, and S2.